OpenStudy (thunder33):

will fan and medal What is the equation of a line that is perpendicular to y=4x+5 and passes through the point (8, 3)? A. y=-1/4x-8 B. y=1/4x-8 C. y=-1/4x+5 D. y=-1/4x+3

Vocaloid (vocaloid):

so, what's the slope of y = 4x + 5?

OpenStudy (welshfella):

Ok a line perpendicular to a line with slope m has slope -1/m

OpenStudy (welshfella):

so if you answer Vocaloid's question you can get the slope of the line perpendicular to given line.

OpenStudy (slayerbodine76):

I will do a perpendicular and a parallel to show you how to do it... First you need to put the equations in y = mx + b form 1. x - 5y = -10 -5y = -x - 10 (multiply the equation by -1) 5y = x + 10 (now divide by 5) y = 1/5x + 2 Perpendicular lines have slopes that are the negative reciprocal. So the slope will be -5. y - y1 = m(x - x1) using slope -5 and points (2,5) y - 5 = -5(x - 2) y - 5 = -5x + 10 y = -5x + 10 + 5 y = -5x + 15 (answer C) 2. 10x - 5y = 8 -5y = -10x + 8 (multiply by -1) 5y = 10x - 8 (divide by 5) y = 2x -8/5 Parallel line have the same slope, so the slope will be 2. y - y1 = m(x - x1) using slope 2 and points (2,4) y - 4 = 2(x - 2) y - 4 = 2x - 4 y = 2x - 4 + 4 y = 2x (answer E)

OpenStudy (thunder33):

so the answer would be C y=1/4x+5

OpenStudy (slayerbodine76):

medal me plz